(e) Simplify a trigonometric expression sin θ cos2 θ − sin θ


1
Expert's answer
2021-12-02T18:55:12-0500
"\sin \theta\cos^2\theta-\sin \theta=-\sin \theta(1-\cos^2\theta)"

"=-\sin \theta\sin^2 \theta=-\sin^3 \theta"